# A Detection of Circumgalactic Dust at Megaparsec Scales with Maximum Likelihood Estimation

### v1

@@ 6 Discussion
Finally, we consider the implications of our choice of dust model. All results presented in this work necessarily assume a form for the dust extinction law $d\mathrm{M}/d\rho$ shown in Equations 4 and 5. In turn, $d\mathrm{M}/d\rho$ necessarily assumes some physical characteristics for dust that govern the wavelength dependence of the extinction. Our adoption of G23 with $R_{V}=3.1$ is tantamount to stating that dust in the CGM is the same as dust in the Milky Way ISM. In reality, dust composition likely differs between the two environments: large dust grains are more likely to survive transport from the ISM to the CGM in hot winds; meanwhile, small grains are likely formed in situ via shattering of larger grains (H. Hirashita, 2024; H. M. Richie et al., 2024). Although clearly not a perfect representation of the CGM, use of a Milky-Way type $d\mathrm{M}/d\rho$ model at least allows for easy comparison with the literature. We plan to evaluate the impact of different dust models on observed $A_{V}$ profiles in future work.
7 Conclusions and Future Directions
@@
In this work, we introduce a new technique for studying circumgalactic dust halos: a maximum-likelihood estimator for dust-induced extinction. Our implementation features up-to-date models for dust-induced reddening (K. D. Gordon et al., 2023) and foreground Galactic extinction (Y.-K. Chiang, 2023). We apply our estimator to two redMaGiC galaxy catalogs with distinct selection functions and perform angular cross-correlation of the resulting extinction with two foreground galaxy samples: a $z<0.2$ subset of WISExSuperCOSMOS galaxies and redMaGiC LRGs at $0.450.15<z<$ 0.45. We detect extinction profiles consistent with megaparsec-scale dust halos in both galaxy samples, with several differences between the two. While the LRG sample displays no extinction in the inner halo ($r\leq 50$ $h^{-1}\,$kpc), the WISExSCOS sample, which includes star-forming galaxies, exhibits a pronounced rise in its extinction profile. At larger scales, the LRG extinction signal remains systematically higher out to 30 $h^{-1}\,$pc, with a noticeable dip at 2 $h^{-1}\,$pc, while the WISExSCOS profile tapers off beyond 1-2 $h^{-1}\,$pc. These patterns are broadly consistent with previous observations of CGM dust and mirror the expected star-forming versus quiescent galaxy populations of the two samples. We perform several tests for systematic error in our pipeline, including substituting star catalogs for both foreground and background catalogs and inverting the redMaGiC LRG samples. These tests indicate no significant biases in our measured extinction profiles, reinforcing the robustness of our main findings.
@@
The authors thank Bryanne McDonough for insightful discussions on the topic of CGM dust and helpful feedback on the manuscript. This work has made use of CosmoHub, developed by PIC (maintained by IFAE and CIEMAT) in collaboration with ICE-CSIC. CosmoHub has received funding from the Spanish government (grant EQC2021-007479-P funded by MCIN/AEI/10.13039/501100011033), the EU NextGeneration/PRTR (PRTR-C17.I1), and the Generalitat de Catalunya. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00.
@@ Appendix A Comparison between SFD and CSFD
Given the acknowledged problems of SFD (K. Yahata et al., 2007; T. Kashiwagi et al., 2013; D. Lenz et al., 2017; Y.-K. Chiang & B. Ménard, 2019), the CSFD formulation is the default Milky Way extinction correction used in this work. It may nevertheless be of interest to compare the profiles obtained with the SFD and the CSFD prescriptions, as shown in Figure 5. While implementing the SFD extinction correction did not significantly alter profile of the high-density redMaGiC background catalog, it noticeably decreased the amplitude of the high-luminosity high-redshift redMaGiC profile at large impact parameters. Put differently, it appears that the SFD prescription added too much extinction to the high-redshift catalog profile at high impact parameters, but with CSFD, the two redMaGiC catalogs produce more consistent $A_{V}$ profiles. On the other hand, compared to the SFD prescription, the CSFD prescription added a significant amount of dust at higher impact parameters for the redMaGIC LRG sample, leading to significant extinction past 1 $h^{-1}\,$pc. Although quite different from the effect for the WISExSuperCOSMOS result, this behavior for LRG is consistent with the idea that SFD overcorrected foreground extinction due to contamination from the LSS. If one assumes that LRGs are a more biased tracer of the LSS than the average WISExSuperCOSMOS galaxy, then it follows that the CSFD prescription adds extinction back in.
@@
For the sake of comparison, we also fit power laws to the SFD-based extinction profiles. The best-fit power laws for the WISExSCOS foreground obtained with the high-density, high-redshift, and combined redMaGiC profiles are given by Equations A1, A2, and A3, respectively:
$A_{V,\,\mathrm{hidens}}(r)$ $=\left(2.0\pm 0.2\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.85\pm 0.02}$ (A1)
$A_{V,\,\mathrm{hiz}}(r)$ $=\left(2.6\pm 0.3\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.73\pm 0.02}$ (A2)
$A_{V,\,\mathrm{stacked}}(r)$ $=\left(2.3\pm 0.5\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.83\pm 0.04}$ (A3)
@@
As above, Equations A1A3 describe the flatter parts of the profiles (50 $h^{-1}\,$kpc $<r<1$ $h^{-1}\,$pc). Finally, the SFD-based LRG profile is best fit by Equation A4:
$A_{V,\,\mathrm{LRG}}(r)=\left(4.2\pm 0.7\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.95\pm 0.04}$ (A4)
@@
The prefactor of Equation A4 matches Equation 11 within error bars, however, the exponents are in a moderately significant 3 $\sigma$ disagreement. Meanwhile, the prefactor and exponent of Equation A3 agrees with Equation 10 within 1 $\sigma$.

### v2

@@ VI Discussion
Finally, we consider the implications of our choice of dust model. All results presented in this work necessarily assume a form for the dust extinction law $d\mathrm{M}/d\rho$ shown in Equations 4 and 5. In turn, $d\mathrm{M}/d\rho$ necessarily assumes some physical characteristics for dust that govern the wavelength dependence of the extinction. Our adoption of G23 with $R_{\mathrm{V}}=3.1$ is tantamount to stating that dust in the CGM is the same as dust in the Milky Way ISM. In reality, dust composition likely differs between the two environments: large dust grains are more likely to survive transport from the ISM to the CGM in hot winds; meanwhile, small grains are likely formed in situ via shattering of larger grains (Hirashita, 2024; Richie et al., 2024). Although clearly not a perfect representation of the CGM, use of a Milky-Way type $d\mathrm{M}/d\rho$ model at least allows for easy comparison with the literature. We plan to evaluate the impact of different dust models on observed $A_{\mathrm{V}}$ profiles in future work.
VII Conclusions and Future Directions
@@
In this work, we introduce a new technique for studying circumgalactic dust halos: a maximum-likelihood estimator for dust-induced extinction. Our implementation features up-to-date models for dust-induced reddening (Gordon et al., 2023) and foreground Galactic extinction (Chiang, 2023). We apply our estimator to two redMaGiC galaxy catalogs with distinct selection functions and perform angular cross-correlation of the resulting extinction with two foreground galaxy samples: a $z<0.2$ subset of WISExSuperCOSMOS galaxies and redMaGiC LRGs at $0.15<z<0.45$. We detect extinction profiles consistent with megaparsec-scale dust halos in both galaxy samples, with several differences between the two. While the LRG sample displays no extinction in the inner halo ($r\leq 50$ $h^{-1}\,$kpc), the WISExSCOS sample, which includes star-forming galaxies, exhibits a pronounced rise in its extinction profile. At larger scales, the LRG extinction signal remains systematically higher out to 30 $h^{-1}\,$Mpc, with a noticeable dip at 2 $h^{-1}\,$Mpc, while the WISExSCOS profile tapers off beyond 1–2 $h^{-1}\,$Mpc. These patterns are broadly consistent with previous observations of CGM dust and mirror the expected star-forming versus quiescent galaxy populations of the two samples. We perform several tests for systematic error in our pipeline, including substituting star catalogs for both foreground and background catalogs and inverting the redMaGiC LRG samples. These tests indicate no significant biases in our measured extinction profiles, reinforcing the robustness of our main findings.
@@
The authors thank Bryanne McDonough for insightful discussions on the topic of CGM dust and helpful feedback on the manuscript. Support for this work was provided by NASA through ADAP grant 18-2ADAP18-0187 (“STARDUST: Searching The hAlo Regions for Dust”). This work has made use of CosmoHub, developed by PIC (maintained by IFAE and CIEMAT) in collaboration with ICE-CSIC. CosmoHub has received funding from the Spanish government (grant EQC2021-007479-P funded by MCIN/AEI/10.13039/501100011033), the EU NextGeneration/PRTR (PRTR-C17.I1), and the Generalitat de Catalunya. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00.
@@ Appendix A Comparison between SFD and CSFD
Given the acknowledged problems of SFD (Yahata et al., 2007; Kashiwagi et al., 2013; Lenz et al., 2017; Chiang and Ménard, 2019), the CSFD formulation is the default Milky Way extinction correction used in this work. It may nevertheless be of interest to compare the profiles obtained with the SFD and the CSFD prescriptions, which is shown in Figure 5. While implementing the SFD extinction correction did not significantly alter the profile of the high-density redMaGiC background catalog, it noticeably decreased the amplitude of the high-luminosity high-redshift redMaGiC profile at large angular separations. Put differently, the SFD prescription seems to undercorrect Galactic extinction in the high-redshift catalog profile at high impact parameters, but with CSFD, the two redMaGiC $A_{\mathrm{V}}$ profiles are more consistent. On the other hand, compared to the SFD prescription, the CSFD prescription results in a significant amount of extinction at high impact parameters for the redMaGIC LRG sample, past 1 $h^{-1}\,$Mpc. Although quite different from the effect for the WISExSuperCOSMOS result, this behavior for LRG is consistent with the idea that SFD overcorrected Galactic extinction due to contamination from the LSS. If one assumes that LRGs are a more biased tracer of the LSS than the average WISExSuperCOSMOS galaxy, then it follows that the CSFD prescription adds extinction back in.
@@
For the sake of comparison, we also fit power laws to the SFD-based extinction profiles. The best-fit power laws for the WISExSCOS foreground obtained with the high-density, high-redshift, and stacked redMaGiC profiles are given by Equations A1, A2, and A3, respectively:
$A_{\mathrm{V,\,hidens}}(r)$ $=\left(2.0\pm 0.2\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.85\pm 0.02}$ (A1)
$A_{\mathrm{V,\,hiz}}(r)$ $=\left(2.6\pm 0.3\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.73\pm 0.02}$ (A2)
$A_{\mathrm{V,\,stacked}}(r)$ $=\left(2.3\pm 0.5\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.83\pm 0.04}$ (A3)
@@
As above, Equations A1A3 describe the flatter parts of the profiles (50 $h^{-1}\,$kpc <r<1<r<1 $h^{-1}\,$Mpc). Finally, the SFD-based LRG profile is best fit by Equation A4:
$A_{\mathrm{V,\,LRG}}(r)=\left(4.2\pm 0.7\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.95\pm 0.04}$ (A4)
@@
The normalization of Equation A4 matches Equation 11 within error bars, however, the exponents are in a moderately significant 3 $\sigma$ disagreement. Meanwhile, the normalization and exponent of Equation A3 agrees with Equation 10 within 1 $\sigma$.
Appendix B Additional Sources of Bias
Leakage of the foreground sample into the background (or vice versa) would introduce a significant bias into our results and change the amplitude of the measured extinction. To minimize the probability of this sample leakage, we imposed relatively wide redshift separations: $\Delta z=$ 0.3–0.4 for the WISExSCOS–redMaGiC cross-correlation and $\Delta z=0.15$ for the LRG sample. These separations exceed the WISExSuperCOSMOS normalized redshift uncertainty ($\sigma_{z}=0.033$) and the redMaGiC uncertainty ($\sigma_{z}=0.017/(1+z)$) by an order of magnitude.
Even so, photometric redshift estimates are subject to catastrophic outliers and wide posterior distributions. To test for sample leakage in the face of these eventualities, we adopt a standard technique from weak gravitational lensing analyses and compute the boost factor profile $B(\theta)=1+w(\theta)$, where $w(\theta)$ is the angular clustering of galaxy positions. Lensing analyses use $B(\theta)$ to account for lens-source clustering when the two populations overlap in redshift (Miyatake et al., 2015; Prat et al., 2022; Sheldon et al., 2004), making it a useful diagnostic for our purposes as well.
Figure 6: Boost factor profiles of redMaGiC background galaxies as a function of angular separation from foreground (WISExSCOS) galaxies.
We calculate $w(\theta)$ using TreeCorr’s implementation of the Landy-Szalay estimator $\xi=(DD-DR-RD+RR)/RR$, which automatically accounts for survey geometry through the inclusion of random catalogs. Uncertainties are estimated from jackknife resampling across 50 sky patches. As shown in the left panel of Figure 6, the resulting $B(\theta)$ for the cross correlation of foreground and background samples is consistent with unity at the $\leq 10^{-3}$ level over all angular scales considered. The corresponding $w(\theta)\simeq 10^{-4}$$10^{-3}$ is one to two orders of magnitude lower than the clustering autocorrelation of redMaGiC galaxies themselves, e.g., Abbott et al. (2022). This result indicates that leakage-driven clustering between the foreground and background samples is negligible, and by extension, that foreground-background sample leakage is unlikely to be a significant problem in our analysis.
An additional, more subtle systematic concerns the role of extragalactic dust in defining background sample membership. If a background galaxy lies near a selection boundary in color or magnitude, dust-induced extinction and reddening can shift it across that boundary, changing which objects enter the extinction measurement. In this scenario, dust does not act solely as an additive color perturbation, but also as a sample selection operator: dusty sightlines may both redden galaxies and preferentially remove the reddest or faintest objects from the background catalog. If this “sample selection reddening” is asymmetric—if more galaxies leave the background sample due to dust-induced reddening than enter it or vice versa—the mean color of the background sample will be shifted, changing the measured amplitude of the extinction signal by an unknown amount.
This effect is inherently difficult to quantify, particularly for catalogs with complex, empirically calibrated selection functions like redMaGiC. The use of multiple background samples (i.e., redMaGiC high-density and high-lum high-$z$) with the same foreground sample (WISExSuperCOSMOS) provides a practical way to bound the magnitude of any such bias. Fitting each redMaGiC profile individually yields:
$A_{\mathrm{V,\,hidens}}(r)=\left(2.5\pm 0.1\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.83\pm 0.01}$ (B1)
$A_{\mathrm{V,\,hiz}}(r)=\left(4.2\pm 0.1\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.95\pm 0.05}$ (B2)
The normalization of the high-density background profile (Equation B1) is approximately 40% lower than that of the high-luminosity, high-redshift background (Equation B2). Because the selection function of the high-luminosity high-redshift redMaGiC sample is similar—but not identical to—the high-density redMaGiC sample, we interpret the difference in normalization as a conservative estimate of the scale at which background sample selection can influence the inferred extinction amplitude. This difference is modest compared to the dynamic range of the measurement and does not alter the qualitative form of the extinction profile. More broadly, this exercise underscores the value of repeating extinction measurements for a given foreground sample with different, independently selected background samples, both as validation and as a means of bounding selection-related systematics.